OpenStudy (anonymous):

In which polygon is the sum of the measures of the interior angles 2160°? A. Nonagon B. 12-gon C. Decagon D. 14-gon

jimthompson5910 (jim_thompson5910):

S = 180(n-2) 2160 = 180(n-2) solve for n to get n = ???